OpenStudy (anonymous):

Solve this equation: X^2(1+4y^2)dx+3x^3ydy=0

OpenStudy (anonymous):

hey, how's it going?

OpenStudy (anonymous):

Still waiting for helpers.

OpenStudy (anonymous):

-3x^3y dy = x^2(1+4y^2) dx

OpenStudy (anonymous):

dy/dx = [x^2(1+4y^2)]/[-3x^3(y)]

OpenStudy (anonymous):

i thought i must group x to 1 side and y to the other side and integrate them?

OpenStudy (anonymous):

oops, you got me, that is what they want you to do

OpenStudy (anonymous):

-3x^3y dy = x^2(1+4y^2) dx

OpenStudy (anonymous):

divide both sides by -3x^3 and 1+4y^2

OpenStudy (anonymous):

(y/(1+4y^2))dy=(x^2/(-3x^3))dx

OpenStudy (anonymous):

you understand?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thank you.